OpenStudy (anonymous):

Find the product of (x + 5)(x − 5). x^2 − 10x + 25 x^2 + 10x + 25 x^2 − 25 x^2 + 25

OpenStudy (anonymous):

answer is third one.

hartnn (hartnn):

\[\large (a+b)(a-b)=a^2-b^2\] here, a=x and b=5 just plug in values. OR you can foil, if you know.

OpenStudy (anonymous):

Thankyou :)